OpenStudy (anonymous):

Verify the identity cos(x + pi/2) = -sin x

zepdrix (zepdrix):

Using our Angle Addition Formula for Cosine:\[\Large\rm \cos\left(x+\frac{\pi}{2}\right)=\cos x \cos \frac{\pi}{2}-\sin x \sin \frac{\pi}{2}\]Understand how to simplify it from there?

OpenStudy (anonymous):

would it be cos x(0) -sin x(1) ?

zepdrix (zepdrix):

Mmm yes very good!

OpenStudy (anonymous):

yayy thank you so much!